OpenStudy (anonymous):

Good Morning! A group of 15 students bought one lottery ticket. The total winnings = $2300. How much did each student win? How much was left over? Please help.

myininaya (myininaya):

how do we know they got the winnings and not someone else?

OpenStudy (anonymous):

you would divide the money among the 15 students: \[\frac{2300}{15}\] which equals $153.33 for each person. Funny thing is, if we give each person $153.33, we get: (153.33)(15) = 2299.95, which means there is 5 cents left over o.O bizzarre!!

OpenStudy (bahrom7893):

15x=2300 x = 2300/15 = 153.33

OpenStudy (bahrom7893):

those five cents are mine!!!!!! xD

OpenStudy (bahrom7893):

for doing your hw.. haha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

It doesn't say that all students won the same amount of money

OpenStudy (anonymous):

does it?

OpenStudy (anonymous):

There are a lot of specifics left out, im just assuming what the question should have said lol

OpenStudy (anonymous):

yea you're right

OpenStudy (bahrom7893):

yea if it didnt then we wouldnt be able to solve it.. and since they bought one lottery ticket.. i guess they'd split up the total of 15.. let us all give each other medals haha

OpenStudy (anonymous):

medals for everyone <.<

myininaya (myininaya):

lol

OpenStudy (anonymous):

oh, and good morning to you too djov :)

OpenStudy (bahrom7893):

lol Mornin everyone!

myininaya (myininaya):

good morning

OpenStudy (anonymous):

I knew it was a divison problem, but wasn't sure if I should round up. Thank so much !!

OpenStudy (anonymous):

Good morning Joemath

OpenStudy (anonymous):

Logging off. Have a good one everybody!